I found I2. I need help finding I1 and I3. Then find the power output of each resistor.
What is the power output of each resistor in this cirouit? Find SIT,/ 2.0012 10.00 V 3.0012 I.) Draw iss loep lli .0012 5.00 the branch, bes sae Value) 10.00 2 3) Repent for 2nd loe 4) Apply locp ruie and unicfiou nult for he
Answer #1

Applying KVL,

10 - 2i1 - 1i2 - 5 - 4i2 - 3i1 = 0

5i1 + 5i2 = 5

i1 + i2 = 1 ..... (i)

10 - 2i1 - 10i3 - 3i1 = 0

5i1 + 10i3 = 10

i1 + 2i3 = 2 ..... (ii)

KCL, i1 - i2 - i3 = 0 ..... (iii)

Solving,

i1 = 0.8 A, i2 = 0.2 A , i3 = 0.6 A ....Ans


power , P = i^2 R

2 ohm: P = 0.8^2 x 2 = 1.28 W

3 ohm: P = 0.8^2 x 3 = 1.92 W

1 ohm: P = 0.2^ x 1 = 0.04 W

4 ohm: P = 0.2^2 x 4 = 0.16 W

10 ohm: P = 0.6^2 x 10 = 3.6 W